### v1

# Cutoff for the Swendsen–Wang dynamics on the complete graph

Abstract
We study the speed of convergence of the Swendsen–Wang (SW) dynamics for the $q$-state ferromagnetic Potts model on the $n$-vertex complete graph, known as the mean-field model. The SW dynamics was introduced as an attractive alternative to the local Glauber dynamics, often offering faster convergence rates to stationarity in a variety of settings. A series of works have characterized the asymptotic behavior of the speed of convergence of the mean-field SW dynamics for all $q\geq 2$ and all values of the inverse temperature parameter $\beta>0$. In particular, it is known that when $\beta>q$ the mixing time of the SW dynamics is $\Theta(\log n)$. We strengthen this result by showing that for all $\beta>q$, there exists a constant $c(\beta,q)>0$ such that the mixing time of the SW dynamics is $c(\beta,q)\log n+\Theta(1)$. This implies that the mean-field SW dynamics exhibits the cutoff phenomenon in this temperature regime, demonstrating that this Markov chain undergoes a sharp transition from “far from stationarity” to “well-mixed” within a narrow $\Theta(1)$ time window. The presence of cutoff is algorithmically significant, as simulating the chain for fewer steps than its mixing time could lead to highly biased samples.
1 Introduction
The $q$-state ferromagnetic Potts model is a classical spin system model central in statistical physics, theoretical computer science, and discrete probability. Given a finite graph $G=(V,E)$, a set of spins or colors $\mathcal{S}=\{1,\dots,q\}$, and an edge interaction parameter $\beta>0$, it is defined as a probability distribution over the configurations in $\Omega=\{1,\dots,q\}^{V}$, which correspond to the spin assignments from $\mathcal{S}$ to the vertices of $G$. Formally, the probability of a configuration $\sigma\in\Omega$ is given by:
$\mu(\sigma)=\frac{1}{Z}\cdot e^{\beta M(\sigma)},$ (1)
where $M(\sigma)$ denotes the number of edges of $G$ with the same spin in both of its endpoints in $\sigma$ (i.e., the monochromatic edges of $G$ in $\sigma$), and $Z$ is the normalizing constant or partition function. The edge interaction parameter $\beta$ is proportional to the inverse temperature in physical applications; the special case $q=2$ corresponds to the classical ferromagnetic Ising model.
The algorithmic problem of efficiently generating approximate samples from  $\mu$ has attracted much attention in theoretical computer science. For the Ising model (i.e., for $q=2$), seminal results from the 1990s [JS93, RW99] provide a polynomial-time approximate sampling algorithm for $\mu$. Additional approximate-sampling algorithms have been discovered since then for the ferromagnetic Ising setting [CGHT16, GJ17], but all have a large running time of at least $O(n^{10})$ where $n=|V|$. For $q\geq 3$, the sampling problem is computationally hard, specifically #BIS-hard [GJ12, GŠVY16], but efficient approximate sampling algorithms are still possible for many interesting families of graphs; see [HPR20, JKP20, CDF+22, CDK24, BG23a].
An algorithm with the potential of providing both faster algorithms for the $q=2$ case and a technology for understanding the feasibility boundary of approximate sampling for $q\geq 3$ is the Swendsen–Wang (SW) dynamics [SW87]. This sophisticated Markov chain exploits the connection between the ferromagnetic Potts model and the random-cluster model [ES88] and was designed to bypass some of the key difficulties associated with sampling from $\mu$ at low temperatures (large $\beta$). In particular, the SW dynamics is conjectured to converge to $\mu$ in $O(n^{1/4})$ steps when $q=2$ on any graph $G$ for any $\beta>0$. This Markov chain was also recently used to the expand the graph families for which approximately sampling can be efficiently done for $q\geq 3$ and $\beta$ large [BG23a].
The SW dynamics transitions from a configuration $\sigma_{t}\in\Omega$ to $\sigma_{t+1}\in\Omega$ as follows:
  1. 1.
    Independently, for every $e=\{u,v\}\in E$ if $\sigma_{t}(u)=\sigma_{t}(v)$ include $e$ in $M_{t}$ with probability $p=1-e^{-\beta}$;
  2. 2.
    Independently, for every connected component $\mathcal{C}$ in $(V,M_{t})$, draw a spin $s\in\{1,\dots,q\}$ uniformly at random, and set $\sigma_{t+1}(v)=s$ for all $v\in\mathcal{C}$.
The SW dynamics is reversible with respect to $\mu$ and thus converges to it [SW87, ES88]. We refer to Step 1 as the percolation step and to Step 2 as the coloring step.
The SW dynamics is a highly non-local Markov chain, updating and potentially changing the spin of each vertex in each step, making its analysis quite difficult in most settings. Still, significant progress has been made in understanding the speed of convergence of this dynamics in various geometries, including the complete graph [CF99, GJ97, LNNP14, GŠV19, GLP19], finite subsets of the $d$-dimensional integer lattice graph ${\mathbb{Z}}^{d}$ [Ull14, BS17, BCSV18, NS19, BCP+21], trees [Hub04, BCŠV22], random graphs [BG21, COGG+23, BG23b], locally-tree-like graphs [BG23a], graphs with sub-exponential growth [BG23a], graphs of bounded [BCV20, BCC+22] or unbounded degree [BZ23], as well as general graphs when $q=2$ [GJ17]. A setting that has attracted particular interest and that will be the focus of this work is the complete graph, also known as the mean-field model.
On the complete graph, it is natural to re-parametrize the model and use $\beta/n$ instead of $\beta$; i.e., $\mu(\sigma)\propto\exp(\frac{\beta}{n}\cdot M(\sigma))$. A detailed connection between the phase transition of the mean-field model, which we describe in detail in Remark 1.2, and the mixing time of the SW dynamics emerged from [LNNP14, GŠV19, GLP19]. The mixing time of a Markov chain is defined as the number of steps until the Markov chain is close in total variation distance to its stationary distribution starting from the worst possible initial configuration, and it is the most standard notion of speed of convergence to stationarity. Formally, if $P$ is the transition matrix of the chain, for $\varepsilon\in(0,1)$,
$\tau_{\rm mix}(\varepsilon)=\max_{\sigma\in\Omega}\min_{t\geq 0}\{\|P^{t}(\sigma,\cdot)-\mu(\cdot)\|_{\textsc{tv}}\leq\varepsilon\},$
and by convention $\tau_{\rm mix}:=\tau_{\rm mix}(1/4)$.
When $q\geq 3$, the mixing time $\tau_{\rm mix}^{\mathrm{SW}}$ of the SW dynamics satisfies:
$\tau_{\rm mix}^{\mathrm{SW}}=\begin{cases}\Theta(1)&\textrm{if}~\beta<\beta_{\textsc{l}};\\ \Theta(n^{1/3})&\textrm{if}~\beta=\beta_{\textsc{l}};\\ e^{\Omega({n})}&\textrm{if}~\beta\in(\beta_{\textsc{l}},\beta_{\textsc{r}});\\ \Theta(\log n)&\textrm{if}~\beta\geq\beta_{\textsc{r}};\end{cases}$ (2)
see [GŠV19, GLP19]. It is known that $\beta_{\textsc{r}}=q$, but $\beta_{\textsc{l}}$ does not have a closed form, and it is given instead implicitly as the root of a certain equation [GŠV19, LL06]. When $q=2$, $\beta_{\textsc{l}}=\beta_{\textsc{r}}=q$ and there is no slow-mixing window; in particular, the mixing time of the mean-field SW dynamics for $q=2$ is $\Theta(1)$ for $\beta<q$, $\Theta(n^{1/4})$ at the critical point $\beta=q$, and $\Theta(\log n)$ for $\beta>q$ [LNNP14].
In this work, we provide sharper results for the mixing time of the SW dynamics in the $\beta>\beta_{\textsc{r}}=q$ regime, further refining our understanding of the SW dynamics in the mean-field model. We show that there exists a constant $c(\beta,q)>0$ such that $\tau_{\rm mix}^{\mathrm{SW}}=c(\beta,q)\log n+\Theta(1)$, which implies that this Markov chain exhibits cutoff. A Markov chain exhibits cutoff if there is a sharp drop in the total variation distance from close to 111 to close 0 in an interval of time of smaller order than $\tau_{\rm mix}$ which is called the cutoff window; i.e., we say that there is cutoff if as $n\to\infty$, for any fixed $\varepsilon\in(0,1)$,
$\tau_{\rm mix}(\varepsilon)-\tau_{\rm mix}(1-\varepsilon)=o(\tau_{\rm mix}).$
We can now state our main result.
Theorem 1.1.
Let $q\geq 2$ and $\beta>\beta_{\textsc{r}}=q$. Then, there exists a constant $c(\beta,q)>0$ such that SW dynamics for the $q$-state mean-field ferromagnetic Potts model exhibits cutoff at mixing time $\tau_{\rm mix}^{\mathrm{SW}}=c(\beta,q)\log n$ with cutoff window $\Theta(1)$; that is, $\tau_{\rm mix}^{\mathrm{SW}}=c(\beta,q)\log n+\Theta(1)$.
We provide the expression for the constant $c(\beta,q)$ in (4).
The only previously known mixing time cutoff result for the SW dynamics was on the integer lattice graph for sufficiently small enough $\beta$ [NS19]; as such, ours is the first such result in the low-temperature SW dynamics. For the much simpler (local) Glauber dynamics, which updates a single randomly chosen vertex in each step, cutoff is known in the mean-field through the high-temperature $\beta<\beta_{\textsc{l}}$ regime [LLP10, CDL+12]. For the special case of the Ising model, cutoff has also been established for the Glauber dynamics for several graph families; see, e.g., [LS13, LS17, Kim22, vdHK+21].
Recall from (2) that the mixing time of the SW dynamics is $\Theta(1)$ when $\beta<\beta_{\textsc{l}}$ and exponentially slow when $\beta\in(\beta_{\textsc{l}},\beta_{\textsc{r}})$, so whether the chain exhibits cutoff is of limited significance in those parameter regimes. It is an interesting question whether the SW dynamics exhibits cutoff at the dynamical critical points $\beta=\beta_{\textsc{l}}$ or $\beta=\beta_{\textsc{r}}$. In the former, mixing is not entirely governed by a “strong drift” towards a typical spin count distribution and relies on the variance of the dynamics as well; as such, it is unlikely that the dynamics exhibits cutoff at $\beta=\beta_{\textsc{l}}$. For $\beta=\beta_{\textsc{r}}$, the SW dynamics exhibits a first-order drift, but the percolation step is critical, which significantly complicates even the asymptotic mixing time analysis.
From an algorithmic perspective, for a Markov chain like the SW dynamics that exhibits cutoff at $c(\beta,q)\log n$, knowing only the asymptotic mixing time, e.g., that $\tau_{\rm mix}^{\mathrm{SW}}=\Theta(\log n)$, is insufficient in practice, since simulating the dynamics for say $\frac{c(\beta,q)}{2}\log n$ steps, is guaranteed to result in a sample with total variation distance close to 111.
We conclude this introduction with some brief remarks about our proof and techniques. We construct a refined, multi-phase coupling that converges to a pair of configurations with roughly the correct (expected) number of vertices in each spin class. (In the mean-field setting when $\beta>\beta_{\textsc{r}}$, under the stationary distribution $\mu$, there is one dominant spin and all other spins are assigned to roughly the same number of vertices.) Our coupling gradually contracts the distance to this type of configuration over two phases: first to a linear distance and then to within $O(\sqrt{n})$ distance. The second phase determines the leading order of the mixing time. Earlier analyses (e.g., [GŠV19]) used simpler arguments that yield the $O(\log n)$ mixing time bound; such arguments treat step deviations “pessimistically” and cannot provide the sharper bound we derive here. Our analysis carefully accounts for how deviations are amortized over time (i.e., we show that early fluctuations do not matter as much).
An innovation of the present argument is the use of a $q\times q$ projection to a partition $\{V_{1},\dots,V_{q}\}$ of $V$ at low temperatures. This projection contains information about the number of vertices assigned each spin in each $V_{i}$ and its convergence is known to imply that of the SW dynamics due to the symmetry of the mean-field model. This idea was used before in [CDL+12, LNNP14, GŠV19] but in the high-temperature $\beta<\beta_{\textsc{l}}$ setting where the percolation steps of the SW dynamics result in sub-critical random (sub)graphs with trivial connected component structures. In the $\beta>\beta_{\textsc{r}}$ regime, we must contend with the existence of a linear-sized connected component and how it is distributed in the partition $\{V_{1},\dots,V_{q}\}$. To address this, we have an additional step in our analysis in which we argue that the fraction of vertices assigned each spin in each $V_{i}$ coalesces to the overall fraction of vertices with that spin in the full configuration before the full configuration reaches a typical spin count (i.e., before the end of second phase of the coupling). This additional guarantee allows us to design the required coupling.
Remark 1.2.
As mentioned earlier, the mixing behavior of the SW dynamics described in (2) is tightly connected to the order-disorder phase transition of the mean-field model which we describe next for completeness. This phase transition occurs at the critical value $\beta=\beta_{\textnormal{{cr}}}(q)$, where $\beta_{\textnormal{{cr}}}(q)=q$ when $q=2$ and
$\beta_{\textnormal{{cr}}}(q)=2\Big{(}\frac{q-1}{q-2}\Big{)}\log(q-1)$
for $q\geq 3$. When $\beta<\beta_{\textnormal{{cr}}}(q)$, the number of vertices assigned each spin is roughly the same with high probability up to lower-order fluctuations. In contrast, when $\beta>\beta_{\textnormal{{cr}}}(q)$ there is a dominant spin in the configuration with high probability. For $q\geq 3$, the model exhibits phase coexistence at the critical threshold $\beta=\beta_{\textnormal{{cr}}}(q)$; this means that at this point, the set of configurations with no dominant spin and the set of configurations with a dominant spin (with all other non-dominant spins assigned to roughly the same number of vertices), contribute each a constant fraction of the probability mass with all other configurations contributing 0 mass (in the limit as $n\rightarrow\infty$). The effects of the phase coexistence phenomenon extends to the window $(\beta_{\textsc{l}},\beta_{\textsc{r}})$ around $\beta_{\textnormal{{cr}}}(q)$; this window is known as the metastability window and coincides with slow-mixing regime for the SW dynamics. For $q=2$, on the other hand, there is no phase coexistence at $\beta_{\textnormal{{cr}}}(q)$ or metastability, so there is no slow mixing.
2 Preliminaries
In this section, we gather a number of standard definitions and results that we will use in our proofs.
2.1 Couplings
A one step coupling of the Markov chain with state space $\Omega$ specifies for every pair of states $(X_{t},Y_{t})$ a probability distribution over $(X_{t+1},Y_{t+1})$ such that the processes $\{X_{t}\}$ and $\{Y_{t}\}$ are each faithful realizations of the chain, and if $X_{t}=Y_{t}$ then $X_{t+1}=Y_{t+1}.$ The coupling time is defined by
$T_{\mathrm{coup}}=\max_{x,y\in\Omega}\min\limits_{t}\{X_{t}=Y_{t}\mid X_{0}=x,Y_{0}=y\}.$
It is a standard fact that if $\Pr(T_{\mathrm{coup}}>t)\leq\varepsilon$, then $\tau_{\rm mix}(\varepsilon)\leq t$; see, e.g., [LP17]. In addition, to bound $\tau_{\rm mix}=\tau_{\rm mix}(1/4)$, it suffices to bound $\Pr(T_{\mathrm{coup}}>t)$ by any constant $\varepsilon<1$, as the bound can then be boosted by repeating the coupling a constant number of times. For establishing cutoff, however, this approach is not feasible, as it would increase the mixing time bound by a multiplicative factor. As such, throughout our proofs, we must track the probability of success of our multi-stage couplings; in particular, our overall coupling must be designed such that $\Pr(T_{\mathrm{coup}}>t)\leq 1/4$.
2.2 Random graph lemmas
Let $\mathcal{G}$ be a random graph distributed as a $\mathcal{G}(n,\lambda/n)$ random graph with $\lambda>0$. Let $L_{i}(\mathcal{G})$ denote the size of the $i$-th largest component of $\mathcal{G}$ (breaking ties arbitrarily) and let $C(v)$ denote the connected component of vertex $v$ in $\mathcal{G}$.
Lemma 2.1 (Lemma 5.7 [LNNP14]).
Let $I$ be the number of isolated vertices in $\mathcal{G}$. For any constant $\lambda>0$, there exists a constant $C>0$ that $\Pr(I\geq Cn)\geq 1-O(n^{-1})$.
For $\lambda>1$ let $\theta(\lambda)$ denote the unique positive solution of the equation
$e^{-\lambda x}=1-x.$ (3)
Lemma 2.2 (Lemma 5.4 [LNNP14]).
For $\lambda>1$, there exist constants $c=c(\lambda)>0$ and $C>0$ such that for any $A>0$
$\Pr(|L_{1}(\mathcal{G})-\theta(\lambda)n|\geq A\sqrt{n})\leq Ce^{-cA^{2}}.$
Let $\mathcal{L}_{1}(\mathcal{G})$ denote the largest component of $\mathcal{G}$ (note that $|\mathcal{L}_{1}(\mathcal{G})|=L_{1}(\mathcal{G})$). Lemma 2.2 and Hoeffding’s inequality, which also holds when sampling without replacement (see [Hoe94]) imply the following.
Lemma 2.3.
Let $U\subseteq V$. For $\lambda>1$, there exist constants $c=c(\lambda)>0$ and $C>0$ such that for any $A>0$
$\Pr(||{\mathcal{L}}_{1}(\mathcal{G})\cap U|-\theta(\lambda)|U||\geq A\sqrt{n})\leq Ce^{-cA^{2}}.$
Finally, from Theorem 1.1 in [JL08] and the reflection principle, we obtain the following.
Lemma 2.4.
  1. 1.
    If $\lambda<1$, there exists $\gamma=\gamma(\lambda)$ such that $\mathrm{E}[\sum_{j\geq 1}L_{j}(\mathcal{G})^{2}]=\gamma n+O(1)$ and $\mathrm{Var}[\sum_{j\geq 1}L_{j}(\mathcal{G})^{2}]=O(n)$.
  2. 2.
    If $\lambda>1$, there exists $\gamma=\gamma(\lambda)$ such that $\mathrm{E}[\sum_{j\geq 2}L_{j}(\mathcal{G})^{2}]=\gamma n+O(1)$ and $\mathrm{Var}[\sum_{j\geq 2}L_{j}(\mathcal{G})^{2}]=O(n)$.
2.3 Drift function
Let us define the function $F:[1/q,1]\rightarrow[0,1]$ as
$F(x)=\frac{1}{q}+\Big{(}1-\frac{1}{q}\Big{)}\theta(\beta x)x,$
with $\theta(\beta x)$ given by (3). This function plays a central role in the analysis of the mixing time of the SW dynamics, as it captures the expected fraction of vertices in the largest color class after one step of the SW dynamics from a configuration with an $x$ fraction of the vertices in its largest color class. As such, we refer to $F$ as the “drift function”. Our proofs will use the following facts about $F$; the first two were previously established in [GŠV19] and we provide a proof of the next two in Appendix A.
Fact 2.5 (Lemma 4 [GŠV19]).
For $\beta>q$, $F$ has a unique fixed point $a(\beta,q)\in(1/q,1]$ and $|F^{\prime}(a(\beta,q))|<1$.
Fact 2.6 (Lemma 8 [GŠV19]).
For $\beta>0$, $F^{\prime}(x)>0$ and $F^{\prime\prime}(x)<0$ for all $x\in(1/\beta,1]$. That is, $F$ is strictly increasing and concave in the interval $(1/\beta,1]$.
Fact 2.7.
There exists $\Delta(q,\beta)>0$ such that $|F^{\prime\prime}(x)|\leq\Delta(q,\beta)$ for all $x\in(0,1]$.
Fact 2.8.
$F^{\prime}(a)=\frac{q-1}{q}\cdot\frac{\theta(a\beta)^{2}}{\theta(a\beta)+(1-\theta(a\beta))\log(1-\theta(a\beta))}\geq\frac{q-1}{q}.$
3 Proof of main result: cutoff when $\beta>q$
[…]
When $\beta>\beta_{\textsc{r}}$, for a configuration $X\sim\mu$, $\alpha(X)$ will concentrate around the $q$-dimensional majority vector
$m:=\Big{(}a,\frac{1-a}{q-1},\dots,\frac{1-a}{q-1}\Big{)}$
or one of its permutations.
We will design a multi-phase coupling, and in the first two phases, we bound the time it takes a copy of the SW dynamics to reach the neighborhood of $m$ from an arbitrary initial configuration. We use the first phase to argue that the SW dynamics reaches constant distance from $m$ and the second phase to further contract the distance to $O(1/\sqrt{n})$; the latter will dominate the order of the mixing time.
Let $X_{0}\in\Omega$ be an arbitrary configuration and $Y_{0}\sim\mu$. We will bound the probability that $\{X_{t}\}$ and $\{Y_{t}\}$ have not coupled after $c(\beta,q)\log n+O(1)$ steps, where
$c(\beta,q)=\Big{(}2\log\Big{(}\frac{1}{F^{\prime}(a)}\Big{)}\Big{)}^{-1}=\frac{1}{2\log\Big{(}\frac{q}{q-1}\cdot\frac{\theta(a\beta)+(1-\theta(a\beta))\log(1-\theta(a\beta))}{\theta(a\beta)^{2}}\Big{)}};$ (4)
see Fact 2.8. From (3) and the fact that $F(a)=a$, we know that $\theta(a\beta)$ is the unique root in (0,1](0,1]( 0 , 1 ] of the equation
$e^{\frac{-\beta y}{q-(q-1)y}}=1-y.$
In the first two phases $\{X_{t}\}_{t\geq 0}$ and $\{Y_{t}\}_{t\geq 0}$ evolve independently. We claim that after $T_{1}=O(1)$ steps $\|\alpha(X_{T_{1}})-m\|_{\infty}\leq\varepsilon$ and $\|\alpha(Y_{T_{1}})-m\|_{\infty}\leq\varepsilon$ for any desired constant $\varepsilon\in(0,1)$.
Lemma 3.1.
For any constants $\varepsilon>0$ and $\delta\in(0,1)$, for all sufficiently large $n$ and any starting state $X_{0}\in\Omega$, there exists $T=O(1)$ such that after $T$ steps, with probability at least $\delta$, the SW dynamics reaches a state $X_{T}$ such that $\|\alpha(X_{T})-m\|_{\infty}\leq\varepsilon$.
A version of this lemma was established in [GŠV19], but we require a slight refinement here so that the probability of success is any constant arbitrarily close to 111.
In the second phase, we claim that after an additional $T_{2}=c(\beta,q)\log n+O(1)$ steps, with both copies of the chain still evolving independently, for any constant $\delta\in(0,1)$ there exists a constant $C=C(\delta)>0$ such that with with probability at least $\delta$:
$\|\alpha(X_{T_{1}+T_{2}})-m\|_{\infty}\leq\frac{C}{\sqrt{n}},~\text{and}~\|\alpha(Y_{T_{1}+T_{2}})-m\|_{\infty}\leq\frac{C}{\sqrt{n}}.$ (5)
This follows from the following lemma.
Lemma 3.2.
Suppose $X_{0}$ is such that $\|\alpha(X_{0})-m\|_{\infty}\leq\varepsilon$ for a sufficiently small constant $\varepsilon>0$ and let $T=c(\beta,q){\log n}+\gamma$ for constant $\gamma>0$. Then, for any constant $\delta\in(0,1)$, there exists a constant $C=C(\gamma,\delta)>0$ such that with probability at least $\delta$, we have $\|\alpha(X_{T})-m\|_{\infty}\leq{C}/{\sqrt{n}}.$
A slightly weaker convergence result was proved in [GŠV19], where it is required that $T=A\log n$ with $A>c(\beta,q)$. This allowed for a technically simpler analysis. To obtain our sharper bound, we devise a refined approach in which we observe that deviations from the expected proportions vector after one step are amortized over the remaining steps of the dynamics, so that early deviations end up contributing less to the overall separation from $m$.
In preparation for the third phase of coupling, we require an additional “stationary-like” property for both copies of the chain. Let $\mathcal{V}=\{V_{1},\dots,V_{q}\}$ be a partition of $V$ where $V_{i}$ contains all vertices of $X_{T_{1}}$ assigned spin $i$. For constant $\lambda\in(0,1)$, we say that $\mathcal{V}$ is a $\lambda$-partition if $\min_{i}|V_{i}|\geq\lambda|V|$. From Lemma 3.1 we know that $\mathcal{V}$ is a $\lambda$-partition with probability at least $\delta$, for any constant $\delta$ arbitrarily close to 111.
Let $\alpha_{1,i}(X_{t})$ denote the fraction of vertices of $V_{i}$ assigned the dominant in $X_{t}$; i.e., there are $\alpha_{1,i}(X_{t})|V_{i}|$ such vertices. The following lemma ensures that during the second phase, in parallel to making progress towards $m$, at some time $T\leq T_{1}+T_{2}$, the fraction of vertices assigned the spin from the dominant spin class in each $V_{i}$ is roughly the same fraction as that from the full configuration $X_{t}$ in $V$. This additional property will allow us to couple the two process in $O(1)$ additional steps in the third phase.
Lemma 3.3.
Let $X_{0}$ be an arbitrary initial configuration and let $\mathcal{V}$ be a $\lambda$-partition of $V$ with constant $\lambda\in(0,1)$. For any $\delta\in(0,1)$, there exist a constant $C>0$ such that with probability at least $\delta$ when $T\geq\frac{\log n}{2\log\big{(}\frac{q}{q-1}\big{)}}+\frac{\log(2/C)}{\log\big{(}\frac{q}{q-1}\big{)}}$, we have $\max_{k\in[q]}|\alpha_{1}(X_{T})-\alpha_{1,k}(X_{T})|\leq{C}/{\sqrt{n}}.$
We note that $(2\log\big{(}\frac{q}{q-1}\big{)})^{-1}<c(\beta,q)$ by Fact (2.8), and so at time $T_{1}+T_{2}$ we reached two configurations that are close to $m$, i.e., satisfy (5), but that also have $\approx\!a|V_{i}|$ vertices assigned the dominant spin in each partition set $V_{i}$.
For the third phase, we consider another lower-dimensional projection of the SW dynamics; in particular, we zoom into a projection of dimension $q\times q$. For $\{X_{t}\}_{t\geq 0}$, let $A_{t}:=\{A_{ij,t}\}_{i,j\in[q]},$ with $A_{ij,t}$ denoting the number of vertices in $V_{i}$ assigned color $j$ in $X_{t}$. Similarly, define $A_{t}^{\prime}$ for $\{Y_{t}\}_{t\geq 0}$.
We have shown that after the first two phases of the coupling, for any constant $\hat{\delta}\in(0,1)$ there is a suitable constant $\hat{C}>0$ such that at time $T=T_{1}+T_{2}$ with probability at least $\hat{\delta}$ the following holds:
  1. A1.
    $\|\alpha(X_{T})-m\|_{\infty}\leq\hat{C}/\sqrt{n}$ and $\|\alpha(Y_{T})-m\|_{\infty}\leq\hat{C}/\sqrt{n}$;
  2. A2.
    $|\alpha_{1,i}(X_{T})-\alpha_{1,i}(Y_{T})|\leq\hat{C}/\sqrt{V_{i}}$ for all $i\in[q]$.
Under these assumptions, we design a coupling such that $A_{T+1}=A_{T+1}^{\prime}$ with probability $\Omega(1)$.
Lemma 3.4.
Let $\lambda\in(0,1)$ be a constant independent of $n$. Let $\mathcal{V}=\{V_{1},...,V_{q}\}$ be a $\lambda$-partition and let $A_{t}$ and $A_{t}^{\prime}$ be two copies of the projection chain such that A1 and A2 above hold. Then, there exists a coupling of $A_{t+1}$ and $A_{t+1}^{\prime}$ such that with probability $\Omega(1)$, it holds that $A_{t+1}=A_{t+1}^{\prime}$.
This lemma implies that we can obtain two configurations such that $A_{T+1}=A_{T+1}^{\prime}$ with probability $\Omega(1)$. Similar couplings were also designed in [LNNP14, GŠV19] but for the high-temperature $\beta<\beta_{\textsc{l}}$ setting where all the spin classes have roughly the same number of vertices and the percolation step is sub-critical in all classes. Here, we have to contend with the presence of a dominant spin class and a super-critical percolation step; this is the reason for the additional step in our analysis (Lemma 3.4) that guarantees that the extra assumption A2 holds after the first two phases.
The final step of our coupling is a simple boosting phase so that the overall probability of success is at least 3/4$\geq$ 3/4≥ 3 / 4.
Lemma 3.5.
Let $X_{0}$ and $Y_{0}$ be configurations such that A1 and A2 hold. Then, there exists a coupling and a constant integer $\ell\geq 0$ such that $\Pr\big{(}\cup_{i=1}^{\ell}\{A_{t+i}=A^{\prime}_{t+i}\}\big{)}\geq\frac{3}{4}.$
In summary, we have shown that there exists a $T=c(\beta,q)\log n+O(1)$ such that $\|A_{T}-A_{T}^{\prime}\|_{\textsc{tv}}\leq 1/4.$ It was already noted in [LNNP14, GŠV19] that the symmetry of the mean-field setting implies that $\|X_{T}-Y_{T}\|=\|A_{T}-A_{T}^{\prime}\|_{\textsc{tv}}$ and thus
$\tau_{\rm mix}^{\mathrm{SW}}\leq T=c(\beta,q)\log n+O(1),$
which completes the proof of the upper bound in Theorem 1.1. The lower bound in this theorem follows straightforwardly from some of the same facts required to establish Lemma 3.2, and it is provided later in Section 4.1.
We comment briefly in the organization of the rest of the paper. The proofs of Lemmas 3.1 and 3.2, as well as that of the mixing time lower bound, are provided in Section 4. Section 5 contains the proof of Lemma 3.3, and Section 6 those of Lemmas 3.4 and 3.5.
4 Convergence to the neighborhood of the majority vector
We first prove Lemma 3.1 which shows convergence to an $O(1)$ ball around $m$ after $T=O(1)$ steps.
Proof Lemma 3.1.
From Lemma 30 in [GŠV19], we have that for any constant $\varepsilon>0$ and any starting state $X_{0}\in\Omega$, there exists $T=O(1)$ such that after $T$ steps, we have that $\|\alpha(X_{T})-m\|_{\infty}\leq\varepsilon$ with probability at least $\gamma=\Omega(1)$. Hence, running the dynamics for $kT$ steps, the probability that we do not reach a configuration $X$ such that $\|\alpha(X)-m\|_{\infty}\leq\varepsilon$ is at most $(1-\gamma)^{k}$, and the result follows. ∎
We proceed with the proof of Lemma 3.2, which ensures that SW dynamics reaches an $O(1/\sqrt{n})$ ball around $m$ after $T=c(\beta,q)\log n+O(1)$ steps. Let
$\kappa(b):=\Big{(}b,\frac{1-b}{q-1},\dots,\frac{1-b}{q-1}\Big{)}.$
Let $F^{k}(x)=F^{k-1}(F(x))$ with $F^{1}=F$; i.e., $F^{k}$ is the result of applying $F$ repeatedly $k$ times. We show first that $\alpha(X_{t})$ is close in infinity norm to $\kappa(F^{t}(\alpha_{1}(X_{0})))$ for $t\geq 0$ and then that, for an appropriate choice of $t$, $\kappa(F^{t}(\alpha_{1}(X_{0})))$ is close to $m$.
Lemma 4.1.
Suppose $X_{0}$ is such that $\|\alpha(X_{0})-m\|_{\infty}<\varepsilon$ for a sufficiently small constant $\varepsilon>0$. Then, for any $t\geq 0$ and any constant $\delta\in(0,1)$, there exists a constant $C(\delta)>0$ such that with probability at least $\delta$
$\|\alpha(X_{t})-\kappa(F^{t}(\alpha_{1}(X_{0})))\|_{\infty}\leq\frac{C}{\sqrt{n}}.$
Lemma 4.2.
Suppose $x_{0}\in(a-\varepsilon,a+\varepsilon)$ where $\varepsilon>0$ is small enough and let $\gamma\in\mathbb{R}$. Then, if $T=c(\beta,q){\log n}+\gamma$, there exists a constant $C(\gamma)>0$ such that
$\frac{|x_{0}-a|C(\gamma)}{\sqrt{n}}-o\Big{(}\frac{1}{\sqrt{n}}\Big{)}\leq|F^{T}(x_{0})-a|\leq\frac{|x_{0}-a|C(\gamma)}{\sqrt{n}}+o\Big{(}\frac{1}{\sqrt{n}}\Big{)}.$
Proof of Lemma 3.2.
Follows from Lemmas 4.1 and 4.2 and the triangle inequality. ∎
We proceed with the proof of Lemma 4.1 for which we will need the following concentration of the fluctuations.
Lemma 4.3.
Suppose $X_{t}$ is such that $\|\alpha(X_{t})-m\|_{\infty}\leq\varepsilon$ for a sufficiently small $\varepsilon>0$. Then, there exists $\gamma_{0}(\beta)$ such that for any $r\geq\gamma_{0}(\beta)$
$\Pr\Big{(}\|\alpha(X_{t+1})-\kappa(F(\alpha_{1}(X_{t})))\|_{\infty}\geq\frac{r}{\sqrt{n}}\Big{)}\leq e^{-\Omega(r)}+O(r^{-2}).$
Proof.
Since $\beta>q$ and $a>1/q$ by Fact 2.5, $a\beta>1$ and $b\beta<1$. By assumption $\|\alpha(X_{t})-m\|_{\infty}<\varepsilon$, so when $\varepsilon$ is small enough, exactly one color class in the percolation step of the SW dynamics is supercritical, and the remaining $q-1$ color classes are subcritical.
Let $C_{1},C_{2},\dots$ denote the connected components after the percolation step from $X_{t}$, sorted in decreasing order by size and let $R=\sum_{i\geq 2}|C_{i}|^{2}$. By Lemma 2.4 there exist constants $\gamma_{0},\gamma_{1}>0$ such that $\mathrm{E}[R]=\gamma_{0}n+O(1)$ and $\mathrm{Var}[R]\leq\gamma_{1}n.$ Then, by Chebyshev’s inequality, for any $\gamma_{2}\geq 1$ and $n$ large we have
$\Pr\left(\left|R-\gamma_{0}n\right|\geq\gamma_{2}\sqrt{\gamma_{1}n}\right)\leq\Pr\left(\left|R-\mathrm{E}[R]\right|\geq\frac{\gamma_{2}}{2}\sqrt{\gamma_{1}n}\right)\leq\frac{4}{\gamma_{2}^{2}}.$ (6)
For $j\geq 2$ and $i\in[q]$, consider the random variables $Q_{j}^{(i)}$ where
$Q_{j}^{(i)}=\begin{cases}|C_{j}|&\text{if }C_{j}\text{ colored }i,~\text{and}\\ 0&\text{otherwise}.\\ \end{cases}$
Let $Z_{i}=\sum_{j\geq 2}Q_{j}^{(i)}$ be the number of vertices colored $i$ in the coloring step. By Hoeffding’s inequality, for any $\gamma_{3}\geq 0$
$\Pr\Big{(}|Z_{i}-E[Z_{i}]|\geq\gamma_{3}\sqrt{n}\mid R\leq\gamma_{0}n+\gamma_{2}\sqrt{\gamma_{1}n}\Big{)}\leq 2\exp\Big{(}\frac{-2\gamma_{3}^{2}}{\gamma_{0}+\gamma_{2}\sqrt{\gamma_{1}/n}}\Big{)},$
and combined with (6) we deduce by a union bound that
$\Pr\big{[}|Z_{i}-E[Z_{i}]|\geq\gamma_{3}\sqrt{n}\big{]}\leq 2\exp\Big{(}\frac{-\gamma_{3}^{2}}{\gamma_{0}+\gamma_{2}\sqrt{\gamma_{1}/n}}\Big{)}+\frac{4}{\gamma_{2}^{2}}.$
Turning our attention to $|C_{1}|$, by Lemma 2.2, there are exists constants $c_{1},c_{2}>0$ such that
$\Pr(||C_{1}|-\theta(\beta\alpha_{1}(X_{t}))\alpha_{1}(X_{t})n|\geq\gamma_{4}\sqrt{n})\leq c_{1}\exp(-c_{2}\gamma_{4}^{2}).$ (7)
If $||C_{1}|-\theta(\beta\alpha_{1}(X_{t}))\alpha_{1}(X_{t})n|<\gamma_{4}\sqrt{n}$, and $|Z_{j}-E[Z_{j}]|<\gamma_{3}\sqrt{n}$ for all $i\in[q]$, we have
$\alpha_{1}(X_{t+1})n$ $\leq\theta(\beta\alpha_{1}(X_{t}))\alpha_{1}(X_{t})n+\frac{n-\theta(\beta\alpha_{1}(X_{t}))\alpha_{1}(X_{t})n}{q}+(\gamma_{3}+\gamma_{4})\sqrt{n}=F(\alpha_{1}(X_{t}))n+(\gamma_{3}+\gamma_{4})\sqrt{n},$
and similarly
$\alpha_{1}(X_{t+1})n\geq F(\alpha_{1}(X_{t}))n-(\gamma_{3}+\gamma_{4})\sqrt{n}.$
In addition, for $j\neq 1$
$\alpha_{j}(X_{t+1})n$ $\leq\frac{n-\theta(\beta\alpha_{1}(X_{t}))\alpha_{1}(X_{t})n}{q}+(\gamma_{3}+\gamma_{4})\sqrt{n}\leq\frac{(1-F(\alpha_{1}(X_{t})))n}{q-1}+(\gamma_{3}+\gamma_{4})\sqrt{n},$
and
$\alpha_{j}(X_{t+1})n\geq\frac{(1-F(\alpha_{1}(X_{t})))n}{q-1}-(\gamma_{3}+\gamma_{4})\sqrt{n}.$
Combining these facts and setting $\gamma_{2}=\gamma_{3}=\gamma_{4}=r/2$, it follows that for $r\geq\gamma_{0}$ and $n$ large enough that
$\Pr\Big{(}\|\alpha(X_{t+1})-\kappa(F(\alpha_{1}(X_{t})))\|_{\infty}\geq\frac{r}{\sqrt{n}}\Big{)}$ $\leq c_{1}\exp(-c_{2}\gamma_{4}^{2})+2\exp\Big{(}\frac{-\gamma_{3}^{2}}{\gamma_{0}+\gamma_{2}\sqrt{\gamma_{1}/n}}\Big{)}+\frac{4}{\gamma_{2}^{2}}$
$\leq\exp(-\Omega(r))+O(r^{-2})$
as claimed. ∎
We are now ready to provide the proof of Lemma 4.1.
Proof of Lemma 4.1.
Let
$\xi_{t}=\|\alpha_{j}(X_{t})-\kappa_{j}(F(\alpha_{1}(X_{t-1})))\|_{\infty};$
this quantity controls the fluctuation of all color classes at time $t$ from their expected value. Then,
$F(\alpha_{1}(X_{t-1}))-\frac{\xi_{t}}{n}\leq\alpha_{1}(X_{t})\leq F(\alpha_{1}(X_{t-1}))+\frac{\xi_{t}}{n}.$
Similarly,
$F(\alpha_{1}(X_{t-2}))-\frac{\xi_{t-1}}{n}\leq\alpha_{1}(X_{t-1})\leq F(\alpha_{1}(X_{t-2}))+\frac{\xi_{t-1}}{n},$ (8)
[…]
$F\Big{(}F(\alpha_{1}(X_{t-2}))-\frac{\xi_{t-1}}{n}\Big{)}-\frac{\xi_{t}}{n}\leq\alpha_{1}(X_{t})\leq F\Big{(}F(\alpha_{1}(X_{t-2}))+\frac{\xi_{t-1}}{n}\Big{)}+\frac{\xi_{t}}{n}.$
By Fact 2.5, $|F^{\prime}(a)|<1$, so for small enough $\varepsilon$, $\max_{|x-a|<\varepsilon}|F^{\prime}(x)|<1$. Also, since $\alpha_{1}(X_{0})\in(a-\varepsilon,a+\varepsilon)$ by assumption, Facts 2.5 and 2.6 imply that so does $F(\alpha_{1}(X_{t-2}))$. Hence, the Taylor expansion of the function $F$ about $F(\alpha_{1}(X_{t-2}))$ yields that for a suitable $\rho\in(0,1)$
$F\Big{(}F(\alpha_{1}(X_{t-2}))+\frac{\xi_{t-1}}{n}\Big{)}\leq F(F(\alpha_{1}(X_{t-2})))+\frac{(1-\rho)\xi_{t-1}}{n},$ (9)
and
$F\Big{(}F(\alpha_{1}(X_{t-2}))-\frac{\xi_{t-1}}{n}\Big{)}\geq F(F(\alpha_{1}(X_{t-2})))-\frac{(1-\rho)\xi_{t-1}}{n}.$ (10)
Therefore,
$F(F(\alpha_{1}(X_{t-2})))-\frac{(1-\rho)\xi_{t-1}}{n}-\frac{\xi_{t}}{n}\leq\alpha_{1}(X_{t})\leq F(F(\alpha_{1}(X_{t-2})))+\frac{(1-\rho)\xi_{t-1}}{n}+\frac{\xi_{t}}{n}.$
Iterating this argument, we deduce that
$F^{t}(\alpha_{1}(X_{0}))-\frac{1}{n}\sum_{i=1}^{t}(1-\rho)^{t-i}\xi_{i}\leq\alpha_{1}(X_{t})\leq F^{t}(\alpha_{1}(X_{0}))+\frac{1}{n}\sum_{i=1}^{t}(1-\rho)^{t-i}\xi_{i},$ (11)
so that
$|\alpha_{1}(X_{t})-F^{t}(\alpha_{1}(X_{0}))|\leq\frac{1}{n}\sum_{i=1}^{t}(1-\rho)^{t-i}\xi_{i}.$ (12)
In a similar fashion, for $j\neq 1$
$\frac{1-F(\alpha_{1}(X_{t-1}))}{q-1}-\frac{\xi_{t}}{n}\leq\alpha_{j}(X_{t})\leq\frac{1-F(\alpha_{1}(X_{t-1}))}{q-1}+\frac{\xi_{t}}{n},$
and from (8) since $F$ is increasing
$\frac{1-F\big{(}F(\alpha_{1}(X_{t-2}))+\frac{\xi_{t-1}}{n}\big{)}}{q-1}-\frac{\xi_{t}}{n}\leq\alpha_{j}(X_{t})\leq\frac{1-F\big{(}F(\alpha_{1}(X_{t-2}))+\frac{\xi_{t-1}}{n}\big{)}}{q-1}+\frac{\xi_{t}}{n}.$
From (9) and (10), we then deduce that
$\frac{1-F^{2}\big{(}\alpha_{1}(X_{t-2}))}{q-1}-\frac{(1-\rho)\xi_{t-1}}{(q-1)n}-\frac{\xi_{t}}{n}\leq\alpha_{j}(X_{t})\leq\frac{1-F^{2}\big{(}\alpha_{1}(X_{t-2}))}{q-1}+\frac{(1-\rho)\xi_{t-1}}{(q-1)n}+\frac{\xi_{t}}{n},$
and iterating as in (11):
$\left|\alpha_{j}(X_{t})-\frac{1-F^{t}\big{(}\alpha_{1}(X_{t-2}))}{q-1}\right|\leq\frac{1}{(q-1)n}\sum_{i=1}^{t}(1-\rho)^{t-i}\xi_{i}.$ (13)
From (12) and (13), it suffices to show that for a suitable constant $C>0$, with probability at least $\delta\in(0,1)$
$\sum_{i=1}^{t}(1-\rho)^{t-i}\xi_{i}\leq C\sqrt{n}.$
For this, let $K>0$ and let us define the sequence
$r_{i}=K\big{(}1-{\rho}/{2}\big{)}^{t-i}\big{(}1+\rho\big{)}^{t-i}$ (14)
By Lemma 4.3, we can choose $K$ large enough such that
$\sum_{i=1}^{t}\Pr(\xi_{i}\geq r_{i}\sqrt{n})\leq\sum_{i=1}^{t}O(r_{i}^{-2})\leq\delta.$
If $\xi_{i}<r_{i}\sqrt{n}$ for all $i\leq t$, then
$\sum_{i=1}^{t}(1-\rho)^{t-i}\xi_{i}\leq K\sqrt{n}\sum_{i=1}^{t}(1-\rho)^{t-i}\big{(}1-{\rho}/{2}\big{)}^{t-i}\big{(}1+\rho\big{)}^{t-i}\leq C\sqrt{n}$
for large enough $C$. It then follows from a union bound that
$\Pr\Big{(}\sum_{i=1}^{t}\xi_{i}(1-\rho)^{t-i}\geq C\sqrt{n}\Big{)}\leq\sum_{i=1}^{t}\Pr(\xi_{i}>r_{i}\sqrt{n})\leq\delta,$
as claimed. ∎
We complete this proof by proving the proof of Lemma 4.2.
Proof of Lemma 4.2.
For ease of notation, set $|F^{\prime}(a)|=\rho<1$. By Fact 2.7, there exists $\Delta\geq|F^{\prime\prime}(x)|$ for all $x\in(a-\varepsilon,a+\varepsilon)$. Let $d_{t}=|a-F^{t}(x_{0})|$. The Taylor expansion of $F$ about $a$ implies that
$\rho d_{t-1}-\frac{\Delta}{2}d_{t-1}^{2}\leq d_{t}\leq\rho d_{t-1}+\frac{\Delta}{2}d_{t-1}^{2}.$ (15)
Iterating,
$\rho^{T}d_{0}-\frac{\Delta}{2\rho}\sum_{i=1}^{T}\rho^{i}d_{T-i}^{2}\leq d_{T}\leq\rho^{T}d_{0}+\frac{\Delta}{2\rho}\sum_{i=1}^{T}\rho^{i}d_{T-i}^{2}.$
Since $d_{t}\leq\varepsilon$ for $t\geq 0$, for $\varepsilon$ small enough, we can deduce from (15), that there exists a constant $\varepsilon_{0}$ such that $\frac{\Delta\varepsilon}{2}\leq\varepsilon_{0}<1-\rho$ such that $d_{t}\leq(\rho+\varepsilon_{0})d_{t-1}$, and thus
$d_{T}\leq\frac{|x_{0}-a|\rho^{\gamma}}{\sqrt{n}}+\frac{\Delta(\rho+\varepsilon_{0})^{2T}}{2\rho}\sum_{i=1}^{T}\Big{(}\frac{\rho}{(\rho+\varepsilon_{0})^{2}}\Big{)}^{i}\leq\frac{\varepsilon\rho^{\gamma}}{\sqrt{n}}+\frac{\Delta(\rho+\varepsilon_{0})^{2T}}{2\rho}\sum_{i=1}^{T}\Big{(}\frac{\rho}{(\rho+\varepsilon_{0})^{2}}\Big{)}^{i}.$
We can choose $\varepsilon_{0}$ so that $\rho<(\rho+\varepsilon_{0})^{2}$. Then, $(\rho+\varepsilon_{0})^{2T}=o(1/\sqrt{n})$ and $\sum_{i=1}^{T}\Big{(}\frac{\rho}{(\rho+\varepsilon_{0})^{2}}\Big{)}^{i}=O(1)$, and since $\rho^{T}=\frac{\rho^{\gamma}}{\sqrt{n}}$, we obtain
$\frac{d_{0}\rho^{\gamma}}{\sqrt{n}}-o\big{(}\frac{1}{\sqrt{n}}\big{)}\leq d_{T}\leq\frac{d_{0}\rho^{\gamma}}{\sqrt{n}}+o\big{(}\frac{1}{\sqrt{n}}\big{)}.$
4.1 Mixing time lower bound
We will use the following result for proving the the lower bound on the mixing time as stated in Theorem 1.1.
Lemma 4.4.
If $X\sim\mu$, then $\|\alpha(X)-m\|_{\infty}\leq A/{\sqrt{n}}$ with probability at least 3/43/43 / 4 for a suitable constant $A>0$.
Very similar results are available in the literature, but we were not able to find this particular variant of the concentration bound, so, for completeness, we provide a proof of Lemma 4.4 in Appendix B.
Proof of Theorem 1.1 (Lower bound).
By Lemma 4.4, if $X\sim\mu$, then $\|\alpha(X)-m\|_{\infty}\leq A/{\sqrt{n}}$ with probability at least 3/43/43 / 4 for a suitable constant $A$. Now, take $X_{0}$ to be a configuration where all vertices are assigned the same spin. By Lemmas 4.1 and 4.2, there exists $\gamma$ such that after $T=c(\beta,q)\log n+\gamma$ steps, with probability 3/4$\geq$ 3/4≥ 3 / 4, we have $\|\alpha(X_{T})-m\|_{\infty}>\frac{A}{\sqrt{n}}$. This implies that $\|X_{T}-\mu\|_{\textsc{tv}}>1/4$ and thus that the SW dynamics has not mixed after $T$ steps. ∎
5 Spin fraction coalescence in partition
We provide in this section the proof of Lemma 3.3. For this, for $k\in[q]$, let $d_{t}^{(k)}=|\alpha_{1}(X_{t})-\alpha_{1,k}(X_{t})|$ and
$\hat{\xi}_{t}^{(k)}=\Big{|}d_{t}^{(k)}-\Big{(}1-\frac{1}{q}\Big{)}\theta(\beta\alpha_{1}(X_{t-1}))d_{t-1}^{(k)}\Big{|}.$
We will need the following concentration bound for $\hat{\xi}_{t}^{(k)}$.
Lemma 5.1.
Let $\{V_{1},\dots,V_{q}\}$ be a $\lambda$-partition of $V$ and suppose $X_{t}$ is such that $\|\alpha(X_{t})-m\|_{\infty}\leq\varepsilon$ for a sufficiently small $\varepsilon>0$. Then, there exists $\gamma_{1}(\beta,q)$ such that for any $r\geq\gamma_{1}(\beta,q)$ and any $k\in[q]$
$\Pr\Big{(}\hat{\xi}_{t+1}^{(k)}\geq\frac{r}{\sqrt{n}}\Big{)}\leq e^{-\Omega(r)}+O(r^{-2}).$
Proof.
Since $\beta>q$ and $a>1/q$ by Fact 2.5, $a\beta>1$ and $b\beta<1$. By assumption $\|\alpha(X_{t})-m\|_{\infty}\leq\varepsilon$, so when $\varepsilon$ is small enough, exactly one color class in the percolation step of the SW dynamics is supercritical, and the remaining $q-1$ are subcritical.
Without loss of generality, let us prove the statement for $V_{1}$ (i.e., for $k=1$), noting that the same argument applies to any other $V_{i}$. Let $C_{1}$, $C_{2}$,… denote the connected components after the percolation step from $X_{t}$, sorted in decreasing order by size, let $R=\sum_{i\geq 2}|C_{i}|^{2}$ and $R_{1}=\sum_{i\geq 2}|C_{i}\cap V_{1}|^{2}$. By Lemma 2.4 there exist constants $\gamma_{0},\gamma_{1}>0$ such that $\mathrm{E}[R]=\gamma_{0}n+O(1)$ and $\mathrm{Var}[R]\leq\gamma_{1}n$. Then, by Chebyshev’s inequality, for any $\gamma_{2}\geq 1$ and $n$ large we have
$\Pr\big{[}R_{1}\geq\gamma_{0}n+\gamma_{2}\sqrt{\gamma_{1}n}\big{]}\leq\Pr\big{[}R\geq\gamma_{0}n+{\gamma_{2}}\sqrt{\gamma_{1}n}\big{]}\leq\Pr\Big{(}R\geq\mathrm{E}[R]+\frac{\gamma_{2}}{2}\sqrt{\gamma_{1}n}\Big{)}\leq\frac{4}{\gamma_{2}^{2}}.$
For $j\geq 2$ and $i\in[q]$, consider the random variables $Q_{j}^{(i)}$ where
$Q_{j}^{(i)}=\begin{cases}|C_{j}\cap V_{1}|&\text{if}~C_{j}\text{ is colored }i;~\text{and}\\ 0&\text{otherwise}.\\ \end{cases}$
Let $Z_{i}=\sum_{j\geq 2}Q_{j}^{(i)}$ be the number of vertices colored $i$ in $V_{1}$. By Hoeffding’s inequality, for any $\gamma_{3}\geq 0$:
$\Pr\Big{(}|Z_{i}-E[Z_{i}]|\geq\gamma_{3}\sqrt{n}\mid R_{1}\leq\gamma_{0}n+\gamma_{2}\sqrt{\gamma_{1}n}\Big{)}\leq 2\exp\Big{(}\frac{-2\gamma_{3}^{2}}{\gamma_{0}+\gamma_{2}\sqrt{\gamma_{1}/n}}\Big{)},$
and combined with (6) we obtain
$\Pr\big{[}|Z_{i}-E[Z_{i}]|\geq\gamma_{3}\sqrt{n}\big{]}\leq 2\exp\Big{(}\frac{-\gamma_{3}^{2}}{\gamma_{0}+\gamma_{2}\sqrt{\gamma_{1}/n}}\Big{)}+\frac{4}{\gamma_{2}^{2}}.$
We consider next $|C_{1}\cap V_{1}|$. By Lemma 2.3, we have that for suitable constants $c_{1},c_{2}>0$ for any $\gamma_{4}>0$
$\Pr\left(\big{|}|C_{1}\cap V_{1}|-\theta(\beta\alpha_{1}(X_{t}))\alpha_{1,1}(X_{t})|V_{1}|\big{|}\geq\gamma_{4}\sqrt{n}\right)\leq c_{1}\exp{(-c_{2}\gamma_{4}^{2})}$
If $||C_{1}\cap V_{1}|-\theta(\beta\alpha_{1}(X_{t}))\alpha_{1,1}(X_{t})|V_{1}|\big{|}<\gamma_{4}\sqrt{n}$, and $|Z_{j}-E[Z_{j}]|<\gamma_{3}\sqrt{n}$ for all $j\in[q]$, we have
$\alpha_{1,1}(X_{t+1})|V_{1}|$ $\leq\theta(\beta\alpha_{1}(X_{t}))\alpha_{1,1}(X_{t})|V_{1}|+\frac{|V_{1}|-\theta(\beta\alpha_{1}(X_{t}))\alpha_{1,1}(X_{t})|V_{1}|}{q}+(\gamma_{3}+\gamma_{4})\sqrt{n}$
$=G(\alpha_{1}(X_{t}),\alpha_{1,1}(X_{t}))|V_{1}|+(\gamma_{3}+\gamma_{4})\sqrt{n},$
where the function $G:[1/q,1]\times[1/q,1]\rightarrow[0,1]$ is defined as
$G(x,y)=\frac{1}{q}+\theta(\beta x)y\Big{(}1-\frac{1}{q}\Big{)}.$
Similarly, we have that
$\alpha_{1,1}(X_{t+1})|V_{1}|\geq G(\alpha_{1}(X_{t}),\alpha_{1,1}(X_{t+1}))n-(\gamma_{3}+\gamma_{4})\sqrt{n},$
and for $j\neq 2$
$\frac{n-\theta(\beta\alpha_{1}(X_{t}))\alpha_{1,1}(X_{t})|V_{1}|}{q}-(\gamma_{3}+\gamma_{4})\sqrt{n}\leq\alpha_{j,1}(X_{t+1})|V_{1}|\leq\frac{n-\theta(\beta\alpha_{1}(X_{t}))\alpha_{1,1}(X_{t})|V_{1}|}{q}+(\gamma_{3}+\gamma_{4})\sqrt{n}.$
Combining these facts and setting $\gamma_{2}=\gamma_{3}=\gamma_{4}=r/4$, it follows that for $r\geq\gamma_{0}$ and $n$ large enough that
$\Pr\Big{(}|\alpha_{1,1}(X_{t+1})-G(\alpha_{1}(X_{t}),\alpha_{1,1}(X_{t}))|\geq\frac{r}{2\sqrt{n}}\Big{)}$ $\leq\Pr\Big{(}|\alpha_{1,1}(X_{t+1})-G(\alpha_{1}(X_{t}),\alpha_{1,1}(X_{t}))|\geq\frac{(\gamma_{3}+\gamma_{4})}{\sqrt{n}}\Big{)}$
$\leq\Pr\Big{(}|\alpha_{1,1}(X_{t+1})-G(\alpha_{1}(X_{t}),\alpha_{1,1}(X_{t}))|\geq\frac{(\gamma_{3}+\gamma_{4})\sqrt{n}}{|V_{1}|}\Big{)}$
$\leq c_{1}\exp(-c_{2}\gamma_{4}^{2})+2\exp\Big{(}\frac{-\gamma_{3}^{2}}{\gamma_{0}+\gamma_{2}\sqrt{\gamma_{1}/n}}\Big{)}+\frac{4}{\gamma_{2}^{2}}$
$\leq e^{-\Omega(r)}+O(r^{-2}).$ (16)
By Lemma 4.3, for large enough $r$
$\Pr\Big{(}|\alpha_{1}(X_{t+1})-F(\alpha_{1}(X_{t}))|\geq\frac{r}{2\sqrt{n}}\Big{)}\leq e^{-\Omega(r)}+O(r^{-2}).$ (17)
Recall that $d_{t+1}^{(1)}=|\alpha_{1}(X_{t+1})-\alpha_{1,1}(X_{t+1})|$ and
$\hat{\xi}_{t+1}=\Big{|}d_{t+1}^{(1)}-\Big{(}1-\frac{1}{q}\Big{)}\theta(\beta\alpha_{1}(X_{t}))d_{t}\Big{|}.$
Since
$F(x)-G(x,y)=\Big{(}1-\frac{1}{q}\Big{)}(x-y)\theta(\beta x),$
it follows from (16) and (17) that
$\Pr\Big{(}\hat{\xi}_{t+1}\geq\frac{r}{\sqrt{n}}\Big{)}\leq e^{-\Omega(r)}+O(r^{-2}),$
as claimed. ∎
We can now provide the proof of Lemma 3.3.
Proof of Lemma 3.3.
Recall that $d_{t}^{(k)}=|\alpha_{1}(X_{t})-\alpha_{1,k}(X_{t})|$ and
$\hat{\xi}_{t}^{(k)}=\Big{|}d_{t}^{(k)}-\Big{(}1-\frac{1}{q}\Big{)}\theta(\beta\alpha_{1}(X_{t-1}))d_{t-1}^{(k)}\Big{|}.$
Then,
$d_{t}^{(k)}$ $\leq\Big{(}1-\frac{1}{q}\Big{)}\theta(\beta\alpha_{1}(X_{t-1}))d_{t-1}^{(k)}+\hat{\xi}_{t}^{(k)}\leq\Big{(}1-\frac{1}{q}\Big{)}d_{t-1}^{(k)}+\hat{\xi}_{t}^{(k)}.$
Iterating, we deduce that
$d_{t}^{(k)}\leq\Big{(}1-\frac{1}{q}\Big{)}^{t}+\sum_{i=1}^{t}\Big{(}1-\frac{1}{q}\Big{)}^{t-i}\hat{\xi}_{i}^{(k)}.$
Let $\hat{K}>0$ and define the sequence $\hat{r}_{i}=\hat{K}\big{(}1-{\varepsilon}/{2}\big{)}^{t-i}\big{(}1+\varepsilon\big{)}^{t-i}.$ By Lemma 5.1, for $\hat{K}$ large enough
$\sum_{i=1}^{t}\Pr\Big{(}\hat{\xi}_{i}^{(k)}\geq\frac{\hat{r}_{i}}{\sqrt{n}}\Big{)}\leq\sum_{i=1}^{t}O(\hat{r}_{i}^{-2})\leq\frac{\delta}{q}.$
If $\hat{\xi}_{i}^{(k)}<\hat{r}_{i}/\sqrt{n}$ for all $i\leq t$, then
$\sum_{i=1}^{t}(1-\varepsilon)^{t-i}\hat{\xi}_{i}^{(k)}\leq\frac{\hat{K}}{\sqrt{n}}\sum_{i=1}^{t}(1-\varepsilon)^{t-i}\big{(}1-{\varepsilon}/{2}\big{)}^{t-i}\big{(}1+\varepsilon\big{)}^{t-i}\leq\frac{C}{2\sqrt{n}},$
for a suitable constant $C>0$. A union bound then implies that
$\Pr\Big{(}\sum_{i=1}^{t}\hat{\xi}_{i}^{(k)}(1-\varepsilon)^{t-i}\geq\frac{C}{2\sqrt{n}}\Big{)}\leq\sum_{i=1}^{t}\Pr\Big{(}\hat{\xi}_{i}^{(k)}>\frac{\hat{r}_{i}}{\sqrt{n}}\Big{)}\leq\frac{\delta}{q}.$
[…]
6 Projection chain
In this section, we prove Lemma 3.4 using the following multinomial coupling from [GŠV19].
Fact 6.1.
[…]
$\Pr(|X-Y|=\sqrt{n}\cdot(c_{1},c_{2},\dots,c_{q}))\geq 1-\gamma.$
We can now provide the proof of Lemma 3.4.
Proof of Lemma 3.4.
We couple $A_{t+1}$ and $A_{t+1}^{\prime}$ as follows. The percolation step is done independently in both copies. In the coloring step, first, the same color is assigned to the largest connected component in both chains, and all other connected components with two or more vertices from both chains are assigned spins uniformly at random independently. In addition, in each $V_{i}$, if one copy has more isolated vertices than the other, the excess of isolated vertices in that copy are assigned spins independently. The coloring of the remaining isolated vertices will be coupled to fix the potential discrepancies created in each $V_{i}$ by these steps as we described next.
For $i,j\in[q]$, let $\hat{a}_{ij}$, $\hat{a}_{ij}^{\prime}$ be the number of vertices assigned color $j$ in $V_{i}$ so far (i.e., excluding the isolated vertices). We claim that there exists a constant $C>0$ such that with probability $\Omega(1)$ it holds for all $i,j\in[q]$ that
$|\hat{a}_{ij}-\hat{a}_{ij}^{\prime}|\leq C\sqrt{|V_{i}|}.$ (18)
To prove this, let $\{C_{k}\}_{k\geq 1}$ be the connected components in the first chain after the percolation step; assume $C_{1}$ is the largest component. Let $C(v)$ be the connected component of vertex $v$ after the percolation step. Note that
$\sum_{k\geq 2}|C_{k}\cap V_{i}|^{2}=\sum_{k\geq 2}\sum_{v\in C_{k}\cap V_{i}}\sum_{w\in C_{k}\cap V_{i}}1=\sum_{k\geq 2}\sum_{v\in C_{k}\cap V_{i}}|C_{k}\cap V_{i}|=\sum_{k\geq 2}\sum_{v\in C_{k}\cap V_{i}}|C(v)\cap V_{i}|\leq\sum_{v\in V_{i}\setminus C_{1}}|C(v)|.$
For $v\not\in C_{1}$, $E[|C(v)|]\leq A_{1}$ for a suitable constant $A_{1}>0$. Hence, taking expectations, it follows from Markov’s inequality that
$\Pr\left(\sum_{k\geq 2}|C_{k}\cap V_{i}|^{2}\geq 100qA_{1}|V_{i}|\right)\leq\frac{1}{100q}.$
A union bound the implies that $\sum_{k\geq 2}|C_{k}\cap V_{i}|^{2}=O(|V_{i}|)$ for all $i\in[q]$ with probability at least 99/10099/10099 / 100.
Let $\hat{n}_{i}$ be the number of vertices in $V_{i}$ excluding those in $C_{1}$ and the isolated vertices. If $C_{1}$ is assigned spin $s$, then conditioning on $\sum_{k\geq 2}|C_{k}\cap V_{i}|^{2}<100qA_{1}|V_{i}|$, by Hoeffding’s inequality, we deduce that for a large enough constant $A_{2}>0$, with probability at least $\frac{99}{100q^{2}}$, for any $i,j\in[q]$ with $j\neq s$, it holds that
$\Big{|}\hat{a}_{ij}-\frac{\hat{n}_{i}}{q}\Big{|}\leq A_{2}\sqrt{|V_{i}|}.$
Similarly, we can get an analogous bound for $\hat{a}_{ij}^{\prime}$, and by a union bound (18) holds for all $i,j\in[q]$ such that $j\neq s$ with probability at least 98/10098/10098 / 100.
By Lemma 2.3 and assumption A2, for a suitable constant $A_{2}>0$, we have for any $i\in[q]$ that
$\Pr\Big{(}||C_{1}\cap V_{i}|-|C_{1}^{\prime}\cap V_{i}||\geq A_{3}\sqrt{n}\Big{)}\leq\frac{1}{100q}.$
Therefore, by a union bound, for all $i\in[q]$
$\Big{|}\hat{a}_{ik}-|C_{1}\cap V_{i}|-\frac{\hat{n}_{k}}{q}\Big{|}$ $\leq(A_{2}+A_{3})\sqrt{n},~\text{and}$
$\Big{|}\hat{a}_{ik}^{\prime}-|C_{1}^{\prime}\cap V_{i}|-\frac{\hat{n}_{k}}{q}\Big{|}$ $\leq(A_{2}+A_{3})\sqrt{n}$
with probability at least 99/10099/10099 / 100, which implies that (18) holds for all $i\in[q]$ for $j=k$ with that probability.
Since $\mathcal{V}$ is a $\lambda$-partition, Lemma 2.1 and a union bound over the partition sets imply that there are at least $\Omega(n)$ unassigned isolated vertices in each $V_{i}$ in both chains with probability $1-O(n^{-1})$. Observe that the distribution of the coloring of the isolated vertices in each $V_{i}$ and each copy is a multinomial distribution. Then, using Fact 6.1, we can couple the coloring of the isolated vertices in each $V_{i}$ to equalize the spin counts and obtain two configurations such that $A_{t+1}=A_{t+1}^{\prime}$ with probability $\Omega(1)$. ∎
We wrap this section with the proof Lemma of 3.5. The idea behind this proof is that we can show that once A1 and A2 are achieved, they are preserved for a constant number of steps with constant probability, providing more attempts for the coupling from Lemma 3.4 to succeed at least once.
To formalize this, let $P_{t}(r)$ be the event at time $t$, we have $\|\alpha(X_{t})-m\|_{\infty}\leq r/\sqrt{n}$, $\max_{k\in[q]}|\alpha_{1}(X_{t})-\alpha_{1,k}|\leq r/\sqrt{n}$, and similarly for $Y_{t}$.
Fact 6.2.
For $\delta\in(0,1)$, there exists $r=r(\delta)>0$ such that $\Pr(P_{t+1}(r)\mid P_{t}(r))\geq\delta$.
Proof.
By assumption, $\|\alpha(X_{t})-m\|_{\infty}\leq r/\sqrt{n}$. Let us assume that $\alpha_{1}(X_{t})\geq a$ (the case when $\alpha_{1}(X_{t})<a$ is analogous). By Fact 2.6, $F(\alpha_{1}(X_{t}))\leq F(a+r/\sqrt{n})$. Moreover, by Fact 2.5 and the mean value theorem, there exists a constant $\eta\in(0,1)$ such that $F(a+r/\sqrt{n})-a=\eta r/\sqrt{n}$ which implies that $(a+r)-F(a+r)=(1-\eta)r/\sqrt{n}$. Hence, in order for $\|\alpha(X_{t+1})-m\|_{\infty}\geq r/\sqrt{n}$, we would need $\|\alpha(X_{t+1})-\kappa(F(\alpha_{1}(X_{t})))\|_{\infty}\geq\frac{(1-\eta)r}{\sqrt{n}}$. From Lemma 4.3, for any large enough constant $r$, we have
$\Pr\Big{(}\|\alpha(X_{t+1})-m\|_{\infty}\geq\frac{r}{\sqrt{n}}\Big{)}\leq\Pr\Big{(}\|\alpha(X_{t+1})-\kappa(F(\alpha_{1}(X_{t})))\|_{\infty}\geq\frac{(1-\eta)r}{\sqrt{n}}\Big{)}\leq O(r^{-2});$
the same holds for $Y_{t+1}$.
Similarly, by assumption, $d_{t}^{(k)}=|\alpha_{1}(X_{t})-\alpha_{1,k}(X_{t})|\leq r/\sqrt{n}$. Then, using Lemma 5.1, for any large enough constant $r$, we have that there exists $\hat{\eta}\in(0,1)$ such that
$\Pr\Big{(}d^{(k)}_{t+1}\geq\frac{r}{\sqrt{n}}\Big{)}\leq\Pr\Big{(}|d^{(k)}_{t+1}-\hat{\eta}d^{(k)}_{t}|\geq\frac{(1-\hat{\eta})r}{\sqrt{n}}\Big{)}\leq O(r^{-2}).$
The same holds for all $k\in[q]$ by a union bound, and for $Y_{t+1}$. The result then follows by another union bound over the four events. ∎
Proof of Lemma 3.5.
For ease of notation, let $P_{t}$ be the event that both assumptions A1 and A2 hold; that is, $P_{t}=P_{t}(\hat{C})$ for $\hat{C}$ sufficiently large. Let $S_{t}$ be the event that $A_{t}=A^{\prime}_{t}$. Then,
$\Pr\big{(}\cup_{i=1}^{\ell}S_{t+i}\big{)}$ $\geq\Pr(\cap_{i=1}^{\ell}P_{t+i-1})\cdot\Pr(\cup_{i=1}^{\ell}S_{t+i}\mid\cap_{i=1}^{\ell}P_{t+i-1})$
$=\Pr(\cap_{i=1}^{\ell}P_{t+i-1})\cdot\big{(}1-\Pr(\cap_{i=1}^{\ell}\neg S_{t+i}\mid\cap_{i=1}^{\ell}P_{t+i-1})\big{)}.$ (19)
For any desired constant $\delta_{2}\in(0,1)$, we have that $\Pr(P_{t+1}|P_{t})\geq\delta_{2}$ by Fact 6.2. Then by a union bound, we have
$\Pr(\cap_{i=1}^{\ell}P_{t+i-1})\geq 1-\ell(1-\delta_{2}).$
(We note that $\delta_{2}$ depends on $\hat{C}$.)
To get an upper bound for that $\Pr(\cap_{i=1}^{\ell}\neg S_{t+i}|\cap_{i=1}^{\ell}P_{t+i-1})$, observe
$\Pr(\cap_{i=1}^{\ell}\neg S_{t+i}|\cap_{i=1}^{\ell}P_{t+i-1})=\Pr(\neg S_{t+\ell}|\cap_{i=1}^{\ell-1}\neg S_{t+i},\cap_{i=1}^{\ell}P_{t+i-1})\cdot\Pr(\cap_{i=1}^{\ell-1}\neg S_{t+i}|\cap_{i=1}^{\ell}P_{t+i-1}).$ (20)
Then, by the Markov property and Lemma 3.4 we have that for some constant $\delta>0$
$\Pr(\neg S_{t+\ell}|\cap_{i=1}^{\ell-1}\neg S_{t+i},\cap_{i=1}^{\ell}P_{t+i-1})=\Pr(\neg S_{t+\ell}|\neg S_{t+\ell-1},P_{t+\ell-1})\leq 1-\delta.$
To bound the other term in (20), note that
$\Pr(\cap_{i=1}^{\ell-1}\neg S_{t+i}\mid\cap_{i=1}^{\ell}P_{t+i-1})=\Pr(\neg S_{t+\ell-1}\mid\cap_{i=1}^{\ell-2}\neg S_{t+i},\cap_{i=1}^{\ell}P_{t+i-1})\cdot\Pr(\cap_{i=1}^{\ell-2}\neg S_{t+i}\mid\cap_{i=1}^{\ell}P_{t+i-1}).$ (21)
Moreover,
$\Pr(\neg S_{t+\ell-1}\mid\cap_{i=1}^{\ell-2}\neg S_{t+i},\cap_{i=1}^{\ell}P_{t+i-1})$ $=\Pr(\neg S_{t+\ell-1}\mid P_{t+\ell-2},P_{t+\ell-1})=\frac{\Pr(\neg S_{t+\ell-1},P_{t+\ell-1}\mid P_{t+\ell-2})}{\Pr(P_{t+\ell-1}\mid P_{t+\ell-2})}$
$\leq\frac{\Pr(\neg S_{t+\ell-1}\mid P_{t+\ell-2})}{\Pr(P_{t+\ell-1}\mid P_{t+\ell-2})}\leq\frac{1-\delta}{\delta_{2}}.$
where the second to last inequality follows from Lemma 3.4 and the fact that $\Pr(P_{t+\ell-1}|P_{t+\ell-2})\geq\delta_{2}$. Generalizing this process, we observe that for integer $j$ with $1\leq j\leq\ell-1$
$\Pr(\cap_{i=1}^{\ell-j}\neg S_{t+i}|\cap_{i=1}^{\ell}P_{t+i-1})=\Pr(\neg S_{t+\ell-j}|\cap_{i=1}^{\ell-j-1}\neg S_{t+i},\cap_{i=1}^{\ell}P_{t+i-1})\cdot\Pr(\cap_{i=1}^{\ell-j-1}\neg S_{t+i}|\cap_{i=1}^{\ell}P_{t+i-1}),$
and
$\Pr(\neg S_{t+\ell-j}\mid\cap_{i=1}^{\ell-j-1}\neg S_{t+i},\cap_{i=1}^{\ell}P_{t+i-1})$ $=\Pr(\neg S_{t+\ell-j}\mid\cap_{k=t+\ell-j-1}^{t+\ell-1}P_{k})$
$=\frac{\Pr(\neg S_{t+\ell-j},\cap_{k=t+\ell-j}^{t+\ell-1}P_{k}\mid P_{t+\ell-j-1})}{\Pr(\cap_{k=t+\ell-j}^{t+\ell-1}P_{k}\mid P_{t+\ell-j-1})}$
$=\frac{\Pr(\neg S_{t+\ell-1},P_{t+\ell-1}\mid\cap_{k=t+\ell-j-1}^{t+\ell-2}P_{k})\cdot\Pr(\cap_{k=t+\ell-j}^{t+\ell-2}P_{k}\mid P_{t+\ell-j-1})}{\Pr(\cap_{k=t+\ell-j}^{t+\ell-2}P_{k}\mid P_{t+\ell-j-1})\Pr(P_{t+\ell-1}\mid\cap_{k=t+\ell-j-1}^{t+\ell-2}P_{k})}$
$=\frac{\Pr(\neg S_{t+\ell-1}\cap P_{t+\ell-1}\mid P_{t+\ell-2})}{\Pr(P_{t+\ell-1}\mid P_{t+\ell-2})}$
$\leq\frac{1-\delta}{\delta_{2}}.$
Hence, we obtain that inductively from (21)
$\Pr(\cap_{i=1}^{\ell-1}\neg S_{t+i}\mid\cap_{i=1}^{\ell}P_{t+i-1})\leq\Big{(}\frac{1-\delta}{\delta_{2}}\Big{)}^{\ell-1},$
and plugging these bounds into (19)
$\Pr(\cup_{i=1}^{\ell}S_{t+i})$ $\geq\Pr(\cap_{i=1}^{\ell}P_{t+i-1})\cdot\Pr(\cup_{i=1}^{\ell}S_{t+i}\mid\cap_{i=1}^{\ell}P_{t+i-1})$
$\geq(1-\ell(1-\delta_{2}))\Big{(}1-(1-\delta)\Big{(}\frac{1-\delta}{\delta_{2}}\Big{)}^{\ell-1}\Big{)}\geq\frac{3}{4}$
for a suitable choice of $\hat{C}$, $\delta_{2}$, and $\ell$. ∎
Acknowledgments
This work transpired from a discussion between the first author, Andreas Galanis, Reza Gheissari, Eric Vigoda, and Daniel Štefankovič during the Workshop on Algorithms and Randomness at Georgia Tech in 2018 in which the high-level idea of the proof was outlined.
Appendix A Proof of facts about the drift function $F$
We include in this appendix the proofs of Facts 2.7 and 2.8 from Section 2.3.
Proof of Fact 2.7.
From direct calculation, it can be checked that
$F^{\prime\prime}(x)=-\frac{(q-1)\,\beta\,\theta(\beta x)\,e^{-x\beta\,\theta(\beta x)}\bigl{(}x\beta\bigl{(}\theta(\beta x)+2e^{-x\beta\,\theta(\beta x)}\bigr{)}-2\bigr{)}}{q\bigl{(}1-x\beta\,e^{-x\beta\,\theta(\beta x)}\bigr{)}^{3}}.$
From (3), we have $x\beta=-\frac{\ln\bigl{(}1-\theta(\beta x)\bigr{)}}{\theta(\beta x)};$ letting let $y=\theta(\beta x)$, we obtain the analogous function:
$G(y)=-\frac{(q-1)\beta y(1-y)\left(-\frac{\ln(1-y)}{y}(2-y)-2\right)}{q\left(1+\frac{\ln(1-y)}{y}(1-y)\right)^{3}}.$
[…]
Proof of Fact 2.8.
From direct calculation of the derivative of $F$, we deduce from (3) that
$F^{\prime}(x)=\frac{(q-1)\theta(\beta x)}{q(1-x\beta e^{-x\beta\theta(\beta x)})}=\frac{q-1}{q}\cdot\frac{\theta(\beta x)^{2}}{\theta(\beta x)+(1-\theta(\beta x))\log(1-\theta(\beta x))}.$
Then, $F^{\prime}(a)\geq\frac{q-1}{q}$ when $\frac{\theta(\beta x)^{2}}{\theta(\beta x)+(1-\theta(\beta x))\log(1-\theta(\beta x))}\geq 1$ which follows from the fact that $\log(1-\theta(\beta x))\leq-\theta(\beta x)$ since $\theta(\beta x)\in(0,1)$. ∎
Appendix B Proof of Lemma 4.4
For completeness, we provide in this appendix a proof of Lemma 4.4. Let $\pi$ be the random-cluster measure on the complete graph $G=(V,E)$ on $n$ vertices with parameters $q$ and $p=1-e^{-\beta}$. That is, for $A\subseteq E$, let
$\pi(A)=\frac{1}{Z_{\textsc{rc}}}p^{|A|}(1-p)^{|E\setminus A|}q^{c(V,A)},$
[…]
We will require the following result for the random-cluster measure, which can be derived straightforwardly from known results in the random graph and random-cluster literature.
Lemma B.1.
Let $A\sim\pi$. Then, there exist constants $\gamma_{0},\gamma_{1}>0$ such that $|L_{1}(A)-\theta n|\leq\gamma_{0}\sqrt{n}$ and $\sum_{i\geq 2}L_{i}(A)^{2}\leq\gamma_{1}n$ with probability at least 0.90.90.9.
Proof.
By Lemma 3.2 in [Bla16], $|L_{1}(A)-\theta n|=O(n^{8/9})$ with probability $1-O(n^{-1})$. Moreover, Lemma 7 in [CDFR00] and Lemma 3.1 in [BGJ96] imply that $L_{2}(A)=O(\log n)$ with probability $1-O(n^{-1})$. Let $\mathcal{B}$ be the event that $|L_{1}(A)-\theta n|=O(n^{8/9})$ and $L_{2}(A)=O(\log n)$ and observe that by a union bound $\Pr(\mathcal{B})\geq 1-O(n^{-1})$.
[…]
$\mathrm{E}[\hat{M}]$ $\leq n^{2}\Pr(R\not\in J)+\sum_{r\in J}\mathrm{E}[M\mid R=r]\Pr(R=r)$
$\leq n^{2}(\Pr(R\not\in J\mid\mathcal{B})+\Pr(\neg\mathcal{B}))+\sum_{r\in J}O(r)\Pr(R=r)$
$\leq O(n)+\mathrm{E}[R]=O(n),$
where in the second inequality we used Lemma 3.1 from [BGJ96] and Lemma 2.4, and in third that by Hoeffding’s inequality $\Pr(R\not\in J\mid\mathcal{B})=O(n^{-2})$. Then, $\mathrm{E}[M]=O(n)$ and by Markov’s inequality there exists a large enough constant $\gamma_{1}>0$ such that $\Pr(M\leq\gamma_{1}n)\geq 0.95$.
Finally, the local limit theorem in Theorem 14 from [LL06] implies that for a suitable constant $\gamma_{0}>0$, we have $|L_{1}(A)-\theta n|\leq\gamma_{0}\sqrt{n}$ with probability at least 0.950.950.95; the result then follows by a union bound. ∎
We now use this lemma to proof Lemma 4.4 using the well-known connection between the random-cluster and the ferromagnetic Potts distributions. In particular, if $A\sim\pi$ and we perform a coloring step of the SW dynamics, the resulting configuration has distribution $\mu$ [ES88].
Proof of Lemma 4.4.
Let $A\sim\pi$ and obtain $X\in\Omega$ by assigning spins uniformly at random from $\{1,\dots,q\}$ to each connected component of $A$ independently; then, $X$ has distribution $\mu$. Let $\mathcal{A}$ be the event hat $|L_{1}(A)-\theta n|\leq\gamma_{0}\sqrt{n}$ and $\sum_{i\geq 2}L_{i}(A)^{2}\leq\gamma_{1}n$. The largest spin class will be the one assigned to the largest component of $A$ and will have expected size $\theta+\frac{1-\theta}{q}=a$; all other colors classes will have expected size $\frac{1-\theta}{q}=\frac{1-a}{q-1}$. Under the assumption that $\sum_{i\geq 2}L_{i}(A)^{2}\leq\gamma_{1}n$, we can prove the concentration of $\alpha(X)$ about $m$ via Hoeffding’s inequality. In particular, for $\ell>0$:
$\Pr\Big{(}\|\alpha(X)-m\|_{\infty}\geq\ell\sqrt{\frac{\gamma_{1}}{n}}\,\Big{|}\,\mathcal{A}\Big{)}\leq 2e^{-2\ell^{2}},$
and for a suitable $\ell>0$ we can bound this probability by 0.10.10.1. Since by Lemma B.1, $\Pr(\mathcal{A})\geq 0.9$, we obtain that
$\Pr\Big{(}\|\alpha(X)-m\|_{\infty}\leq\ell\sqrt{\frac{\gamma_{1}}{n}}\Big{)}\geq 3/4$
as desired. ∎